40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 27 03:56:52 2015, ericq, Update, LSC, DRFPMI work 
    Reply  Thu Oct 29 03:25:49 2015, ericq, Update, LSC, DRFPMI work DRFPMI.pdfDRMIarms.pdf
       Reply  Tue Nov 3 03:12:46 2015, ericq, Update, LSC, DRFPMI work 
Message ID: 11722     Entry time: Thu Oct 29 03:25:49 2015     In reply to: 11718     Reply to this: 11726
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI work 

[ericq, Gautam]

The length of DRFPMI lock did not increase much tonight, but we got a ~80 second sensing matrix measurement, and got the CARM bandwidth up to 10k with two boosts on. 

NB: I did not measure the CARM loop gain at its excitation frequency, so the plotted sensing element is supressed by the CARM loop. However, this is still useful for gauging the size of the PRCL signal vs. the residual CARM fluctuations. The excitations are fairly closely spaced between 309 and 316 Hz. 

For comparison, I'm also re-plotting the DRMI sensing measurement from a few weeks back taken at CARM offset of -4. We can see some change in the PRCL sensing, likely due to the CARM-coupled path. MICH/PRCL sadly looks pretty degenerate, but REFL55 looks more reasonable. 

 

I think the main limitation tonight was SRC stability. Even before bringing CARM to zero offset, we would see occasional sharp dives in AS110 power. One lockloss happened soon after such an occurance, but I checked the values, and it was not sufficient to trigger the Schmitt trigger down; instead it may have been a real optical loss of signal. The SRCL OLTF looks sensible. 

Random notes:

  • Aux X laser was glitching yet again, twiddled laser current to 1.90A from the 1.95A that I twiddled it to on Monday from the nominal 2.0A. 
  • When aligning the PRMI, I saw both ITMs' oplevs shift by a few urad in both pitch and yaw when engaing/breaking the lock, but this was not repeatable.
  • I reduced the AS110 whitening gain by 9dB, since the DC values were a few thousands, and I wanted to make sure there were no stray ADC saturations. This didn't change lock stability though. 
Attachment 1: DRFPMI.pdf  97 kB  | Hide | Hide all
DRFPMI.pdf
Attachment 2: DRMIarms.pdf  80 kB  | Hide | Hide all
DRMIarms.pdf
ELOG V3.1.3-